Hello, I need help solving this linear algebra problem.

1. Let L be the set of all linear transforms from R3 to R2.

(a) Verify that L is a vector space.

(b) Determine the dimension of L and give a basis for L.
Answer #1

1. - 10 1 L = FT: T IR3 (Rt is lineur map Let I,UEL and der (T+u) (x + y) = (a+y) + u(x+y) Tel + Tryl + vajtu yg = 5 x + 1 Ta4 (T+ U) (V) = TN + UU to VER Ww & Tui = = (U+T)(w for allvar I Ttu= UIT 0 (T+O) U = Tus+ 0 = Touto L = Tow for all VG (R3 7T for all ut (R² VEIRS 7 Copy atRT) 9 (1.T) (W = 1.Tu = T (1.1) - Tcr for all a F-T 0 L. (I + U) () = x (T(v) + U) - DIU + avT = A where = { liger, egy is standard basis of R3 and r = 40,0), (011)} IT) is 3x3 matinx- Je Mexa Cor). Since, s is buses ohet TE l be any linear map * ProF: 01) taan (T) + as [is] p + de 178 - I G Span & Tin а ; Т. 2 р T т 73 km2 2 4:20 { Tig o 3